OpenStudy (anonymous):

Let f(x) = x + 2 and g(x) = x2 - 6x + 3. Find g(f(x))

OpenStudy (theviper):

@ganeshie8 plz help :)

OpenStudy (moongazer):

g(f(x)) = g(x + 2) = (x + 2)^2 - 6(x + 2) + 3

OpenStudy (moongazer):

do you know what's next?

OpenStudy (anonymous):

no :(

OpenStudy (anonymous):

do you distribute the ^2 and the -6 into each parenthesis?

OpenStudy (moongazer):

(x + 2)^2 = (x + 2)(x + 2) - 6(x + 2) = -6(x) + -6(2)

OpenStudy (anonymous):

so would the final answer be x^2 - 2x - 5?

OpenStudy (moongazer):

yup :)

OpenStudy (anonymous):

THANKS!

OpenStudy (moongazer):

this might help you http://www.khanacademy.org/math/algebra/algebra-functions/v/functions-part-2